a harmonic oscillator has a mass of 0.20kg and unknown spring constant, if the period is T = 20s, find k
1
Expert's answer
2019-05-17T11:52:22-0400

The period is given by the formula:


"T=2\pi \sqrt{\frac{m}{k}}"

"k=m\left(\frac{2\pi}{T}\right)^2"

"k=0.2\left(\frac{2\pi}{20}\right)^2=0.020\frac{N}{m}"